Pembroke Co. wants to issue new 19-year bonds for some much-needed expansion projects. The company currently has 10 percent coupon bonds on the market that sell for $1,050, make semiannual payments, and mature in 19 years. What coupon rate should the company set on its new bonds if it wants them to sell at par? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) Coupon rate
Answer #1

We have to calculate required return of the bond, which can be calculated using the excel function RATE:

Total periods shall be 19*2 = 38

PMT shall be equal to 10%/2 * 1000 = 50

PV = 1050

FV = 1000

=RATE(38,50,-1050,1000) = 4.714%

As the rate is semianually, it will be multiplied by 2, hence the answer will be: 9.43%